# Describing Globally Distributed Software Architectures for Tax Compliance

### v2

@@ Abstract.
Background: The company-internal reuse of software components owned by organizational units in different countries is taxable.
@@
Objective: In this article, we introduce the concerns of tax authorities as stakeholders and investigate how software companies can describe their globally distributed software architectures to tax authorities.
@@
Method: In an experimental simulation, we (1) develop a viewpoint that frames the concerns of tax authorities, (2) create a view of a large-scale, globally distributed microservice architecture from a multinational enterprise, and (3) evaluate the resulting software architecture description with a panel of four tax experts.
@@
Results: The panel found our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. The architecture description reveals that almost 70% of all reuse relationships between the 2560 microservices from our case company are cross-border and, therefore, taxable. However, unclear jurisdictions of owners and potentially insufficient definitions of code ownership and software component introduce significant noise to the view that limits the usefulness and explanatory power of our software architecture description.
Conclusion: Although our software architecture description already provides a solid foundation and reveals the importance of tax compliance in software architectures, we stumbled over several fundamental open questions, forming new frontiers in software engineering.
@@
software architecture, taxation, globally distributed software development
copyright: acmcopyrightisbn: journal: tosemccs: Software and its engineering Software organization and properties
“We don’t do this because it is easy but because we thought it would be easy.”
—Anonymous
1. Introduction
Code is not tied to a physical, geographical location due to its intangible nature. The code owner of a software component, however, is. Code ownership refers to assigning responsibility and accountability for a particular piece of code or software component to an individual, a team, or an organizational unit within an organization.
@@ 1. Introduction
The geographical location of those software component owners has a legal implication that is too often underestimated: Using software components with owners in different countries is taxable as it constitutes an often implicit form of licensing (Dorner et al., 5555). To comply with the international standard of taxation (OECD, 2022), multinational enterprises must be able to report a geographical perspective on their software architecture. However, there is no architectural viewpoint that frames the concerns of tax authorities as stakeholders towards a globally distributed software architecture, which could serve as the foundation for a software architecture description.
@@
In this article, we investigate the concerns of tax authorities as stakeholders of globally distributed software architectures and seek an answer to the research question:
@@
How can multinational enterprises describe their globally distributed software architectures to tax authorities?
We address this research question in an experimental simulation with a multinational enterprise aiming to report an architecture description of its globally distributed software architecture. Instead of all involved tax authorities, we build a panel of four tax experts evaluating the usefulness and completeness of the two core elements of the software architecture description, namely the architecture viewpoint and the architecture view. This experimental design enables us to focus on evaluating emerging technical, procedural, and legal issues in practice while reporting software architecture descriptions to tax authorities without any potential legal implications and consequences—which is crucial in this sensitive context.
The main contributions of this study are the following:
  • An introduction to the concerns of tax authorities as stakeholders of globally distributed software architectures
  • A novel software architecture viewpoint specifying the conventions for constructing and using a view for the concerns of tax authorities towards geographically distributed software architectures
  • An empirical view of an industrial globally distributed microservice architecture using our novel viewpoint
  • An evaluation of the software architecture description for tax compliance by a panel of tax experts
Throughout this paper, we use the terminology and definitions of ISO/IEC/IEEE 42010:2022 (ISO/IEC/IEEE42010:2022, 2022). In particular, we distinguish between architecture description, architecture view, and architecture viewpoint: An architecture description includes one or more architecture views. An architecture view (or simply, view) addresses one or more of the concerns held by the system’s stakeholders. An architecture view expresses the architecture of the system of interest in accordance with an architecture viewpoint (or simply, viewpoint). There are two aspects to a viewpoint: the concerns it frames for stakeholders and the conventions it establishes on views. Viewpoint conventions can include languages, notations, model kinds, design rules, and/or modelling methods, analysis techniques, and other operations on views.
@@
The remainder of the article is structured as follows: After establishing the background and related work in Section 2 and discussing our study design as an experimental simulation in Section 3, we define the viewpoint (Section 4.2) and how we use it to generate a view for the concern of tax authorities (Section 4.3). In Section 5, we present and discuss the evaluation results for our novel viewpoint (Section 5.1 and the view (Section 4.3) in detail. We conclude the paper with a summary of our findings and the implications for research and practice before outlining future work. Figure 1 provides an overview of our study.
[…]
Figure 1. An overview of the structure of our study as experimental simulation and its two components: (1) creating a software architecture description with a viewpoint derived from the concerns of tax authorities and a view of an empirical software architecture using the viewpoint and (2) evaluating the software architecture description with a panel of tax experts.
2. Related work
@@
This article connects taxation with software architecture distributed across multiple countries. Therefore, we provide related work on globally distributed software development and the emerging field of taxation in software engineering in the following subsections.
2.1. Globally distributed software development
@@
Further, strict code ownership is also found to introduce communication overhead and delays (Clerc et al., 2007). For these reasons, and due to the increased use of agile methods, there has been an increasing interest in collaboratively developing software products. This includes a shift from modularization based on geographically distributed locations towards componentization through microservices (Newman, 2021) on a team level, as well as a trend towards establishing collective ownership of source code and other documents to allow everyone in the project to work on any model or artifact (Clerc et al., 2007). Collaborative work includes keeping track of which team owns and does what (e.g., through Scrum of Scrum fora or organizing cross-unit demos), submitting contributions to other’s components and reviewing each other’s code (Dorner et al., 5555; Šmite et al., 2023), and fostering reuse of artifacts to reduce redundant work.
2.2. Taxation in software engineering
@@
Taxation in software industry has been debated for many decades (OECD, 2015). The problem with taxing the final result of software engineering, the software product or service, for example, has shown to be challenging to tackle and is still subject to ongoing and broad discussion (Olbert and Spengel, 2017). This debate has also reached software engineering, the construction of the software product; at first in the context of inner source, use of open source software development practices and the establishment of open source-like communities within an organization (Capraro, 2020). To address the taxation challenges in inner source software development, Buchner and Riehle proposed a cost-based approach (Buchner and Riehle, 2022). However, Treidler et al. showed that a cost-based approach is not applicable since inner source does not reflect the function and risk profile, the starting point for any transfer pricing discussion, for a cost-based approach (Treidler et al., 2024).
@@
Dorner et al. identified the use of software components owned by teams or individuals who represent separate geographically distributed subsidiaries of the same enterprise as type of licensing—a key intangible in software engineering that is taxable (Dorner et al., 5555). For an in-depth discussion on the concerns of tax authorities on software architecture, we refer the reader to Section 4.1 in which we discuss a simplified example of a multinational enterprise developing software and how tax authorities become a stakeholder to a globally distributed software architecture.
3. Study design
We designed our study as an experimental simulation (Stol and Fitzgerald, 2018). In such a setup, the behaviors of actors that a researcher aims to observe and measure are natural and rooted in their professional experience and expertise, but the setting in which these behaviors occur has been specifically created for the purposes of the study. That is, without the study, the setting would not exist. Thus, the term “simulation” refers in this context to the artificially created setting that aims to recreate a concrete type of setting in which the “experiment” (i.e., observed behavior) takes place (Stol and Fitzgerald, 2018).
In our study, we simulate how software companies can describe their globally distributed software architectures for tax compliance. The simulation consists of two parts: (1) creating a view using our viewpoint at a multinational enterprise and (2) evaluating by a panel of tax experts in which we observe and measure the usefulness and completeness of an architecture description for tax compliance.
On the one hand, the behaviors of actors we aim to observe and measure are all natural: (1) the multinational enterprise fulfilling its reporting obligations regarding its globally distributed software architecture to comply with international taxation standards, (2) the tax advisors representing and consulting companies and creating a line of reasoning, and (3) tax auditors from tax authorities scrutinizing the line of reasoning. We instructed all participants to imagine it to be in the context of a tax audit.
On the other hand, the setting is specifically created for the purpose of this study, also as a means to ensure anonymity and immunity. Obviously, evaluating our proposed architecture description in a real tax audit (i.e., a natural, real setting) is not possible: Tax audits are highly sensitive and, if not brought to court, confidential. Our setting must, therefore, provide a safe space for all participants without any legal implications for the company or the tax authorities.
However, our setting is not plucked out of thin air but emulates a tax audit as it involves all relevant parties: The panel of four tax experts consists of two tax auditors and two tax advisors representing both opposing perspectives, the company perspective as the provider and the tax authority perspective as the recipient of such documentation. Their expertise and experiences will give us a first evaluation of our preparatory steps and the resulting software architecture description of globally distributed software architectures for tax compliance.
Although this evaluation is a crucial pillar of our study, our study goes beyond a traditional judgment study as it also describes the practical issues for creating and using a view based on our novel viewpoint in an empirical case. We discuss our approach to the limitations in Section 7.
Last but not least, we would like to emphasize that an in-depth legal and economic discussion and evaluation of our work is beyond our reach. We further discuss this limitation and its implications for research in Section 8.
3.1. Data collection
All involved parties want to remain anonymous, including our case company, which was used as a source for creating our view. Therefore, we are not able to describe the case and the data collection in detail. However, we may well argue that our case company is exemplary for a multinational enterprise developing software.
In brief, to create the view, we extracted the software architecture and the team information from an internal system that extracts the component structure of the system from the code repositories on a daily basis.
To evaluate the usefulness and completeness of our software architecture description for its intended purpose, we conducted semi-structured interviews with four tax experts. The main goal of the evaluation is to seek generalizability over the responses rather than generalizability to a population of actors (Stol and Fitzgerald, 2018). We ensured privacy and confidentiality for all actors (i.e., the involved companies, tax authorities, and tax advisors) by conducting the interviews separately and removing all company-specific information. We also altered the geographical distribution of the subsidiaries to make our company not identifiable in any sense on the one hand but also to give the reader a better impression of globally distributed software architecture on the other hand.
The interview guidelines are publicly available [^1] https://zenodo.org/records/10214942. The group of four tax experts represents the two opposing parties in a tax audit: internal or external tax advisors representing the company and tax auditors representing the tax authorities.
We conducted two pilot interviews to address didactical challenges since the interviews were conducted with tax experts with no software engineering background. A pair of a co-author with a taxation background and a co-author with a software engineering background conducted the interviews with the tax experts virtually between 2023-10-24 and 2023-11-03. Table 1 provides an overview of our interviewees and their backgrounds. We recorded all interview sessions.
@@ A Tax auditor
Former tax auditor, majoring in international taxation, OECD advisor for transfer pricing from 2016 to 2018, currently not involved in transfer pricing
K Tax advisor
@@
Transfer pricing advisor, previously transfer pricing expert at one of the Big Four accounting firms for 20 years
M Tax auditor
@@
Tax auditor and key expert on transfer pricing for other tax auditors for more than eight years
R Tax advisor
@@
Managing director at a transfer price consulting firm; tax inspector before, 35 years of experience in international taxation
Table 1. An overview of the interviewees and their backgrounds.
3.2. Data analysis
@@
As our interview guidelines show, large parts of the interviews were dedicated to providing contextual software engineering understanding for the tax experts. Therefore, we translated and paraphrased only the answers from the interviewees. While one researcher of the interview team analyzed the recorded interviews using thematic analysis (Clarke and Braun, 2017), the second one ensured the quality of the findings and mappings. After the coding, we again reflected on the findings.
4. A software architecture description for tax compliance
@@
In the following subsections, we provide a gentle introduction to international taxation for software architects and specify the set of conventions for creating, interpreting, and using an architecture view to frame the concern of tax authorities.
4.1. International taxation standards as concern and tax authorities as stakeholder
@@
Consider the fictional devnullsoft Group, a multinational enterprise that develops and sells a software product. The software is split up into 18 components. Each component has an owning team that is responsible and accountable for its components. The teams are assigned to one of the legal entities of the enterprise in three countries depending on the location of the team members: devnullsoft GmbH in Germany, devnullsoft AB in Sweden, and devnullsoft Ltd. in the UK. In Figure 2, we provide a schematic overview of a simplified globally distributed software architecture with owning teams assigned to three different subsidiaries in Germany, Sweden, and the UK. Using code components across different jurisdictions is taxable.
@@
Figure 2. A schematic overview of a globally distributed software architecture developed by the fictional devnullsoft Group, a multinational software enterprise with three subsidiaries developing software. Using code components across different jurisdictions is taxable.
@@
All three questions must be answered in the context of software architecture.
4.2. Architecture viewpoint
@@
4.2.1. Defining component and component owner
@@
4.2.2. Decomposing the component structure of the software product
@@
4.2.3. Identifying component owners
@@
4.2.4. Identifying jurisdictions of component owners

### v3

@@ Abstract
Background: The company-internal reuse of software components owned by organizational units in different countries constitutes an implicit licensing across borders, which is taxable. This makes tax authorities a less known stakeholder in software architectures.
@@
Objective: Therefore, we investigate how software companies can describe the implicit license structure of their globally distributed software architectures to tax authorities.
@@
Method: We develop a viewpoint that frames the concerns of tax authorities, use this viewpoint to construct a view of a large-scale microservice architecture of a multinational enterprise, and evaluate the resulting software architecture description with a panel of four tax experts.
@@
Results: The panel found our proposed architectural viewpoint properly and sufficiently frames the concerns of taxation stakeholders. However, unclear jurisdictions of owners and potentially insufficient definitions of code ownership and software component introduce significant noise to the view that limits the usefulness and explanatory power of our software architecture description.
Conclusion: While our software architecture description provides a solid foundation, we believe it only represents the tip of the iceberg. Future research is necessary to pave the way for advancements in tax compliance within software engineering.
Keywords:
@@
software architecture, taxation, globally distributed software development, microservices, tax compliance
1 Introduction
@@
Code is not tied to a physical, geographical location due to its intangible nature. The code owner of a software component, however, is. Code ownership refers to assigning responsibility and accountability for a particular piece of code or software component to an individual, a team, or an organizational unit within an organization. However, the geographical location of those software component owners has a legal implication that is too often underestimated: Using software components with owners in different countries is taxable as it constitutes from a taxation perspective an—often implicit—form of licensing (Dorner et al., 2024; OECD, 2022). To comply with the international standard of taxation (OECD, 2022), multinational enterprises must be able to report this implicit license structure of their software architectures from a geographical perspective. However, there is no architectural viewpoint that frames these concerns of tax authorities as stakeholders towards a globally distributed software architecture.
@@
In this study, we aim to understand the concerns of tax authorities as stakeholders towards software architectures in multinational enterprises and seek an answer to the research question:
@@
How can multinational enterprises describe the implicit license structure of their globally distributed software architectures to tax authorities?
Therefore, we developed a viewpoint that frames the concerns of tax authorities, use this viewpoint to create a view of a large-scale microservice software architecture of a multinational enterprise, and evaluate the resulting software architecture description with a panel of tax experts.
The main contribution of this study is the presentation and evaluation of a novel software architecture description for tax compliance.
Throughout this paper, we use the terminology and definitions of ISO/IEC/IEEE 42010:2022 . In particular, we distinguish between architecture description, architecture view, and architecture viewpoint: An architecture description is the work product used to express software architectures and is a tangible representation of information provided to the stakeholders. An architecture view (or simply, view) is an information part comprising a portion of an architecture description. An architecture view expresses the architecture of the system of interest in accordance with an architecture viewpoint (or simply, viewpoint). A viewpoint is a set of conventions for the creation, interpretation, and use of a view. The software architecture description under evaluation has only one view.
@@
The remainder of the article is structured as follows: After establishing the background and related work in Section 2 and discussing our study design as a judgment study in Section 3, we define the viewpoint (Section 4.2) and how we use it to generate a view for the concern of tax authorities (Section 4.3). In Section 5, we present the evaluation results for our novel viewpoint (Section 5.1, the view (Section 5.2), and the overall software architecture description (Section 5.3) in detail. After critically reflecting on the evaluation results in Section 6 and the limitations of our study in Section 7, we conclude the paper with a summary of our findings and the implications for research and practice before outlining future work in Section 8.
At this point, we would like to emphasize that we focus on a software-engineering perspective on tax compliance. An in-depth legal and economic discussion and evaluation of our work is beyond our reach.
2 Related Work
@@
Since this article connects taxation with software architecture distributed across multiple countries, we provide related work on globally distributed software development and the emerging field of taxation in software engineering in the following subsections.
2.1 Globally Distributed Software Development
@@
Further, strict code ownership is also found to introduce communication overhead and delays (Clerc et al., 2007). For these reasons, and due to the increased use of agile methods, there has been an increasing interest in collaboratively developing software products. This includes a shift from modularization based on geographically distributed locations towards componentization through microservices (Newman, 2021) on a team level, as well as a trend towards establishing collective ownership of source code and other documents to allow everyone in the project to work on any model or artifact (Clerc et al., 2007). Collaborative work includes keeping track of which team owns and does what (e.g., through Scrum of Scrum fora or organizing cross-unit demos), submitting contributions to other’s components and reviewing each other’s code (Dorner et al., 2024; Šmite et al., 2023), and fostering reuse of artifacts to reduce redundant work.
2.2 Taxation in Software Engineering
@@
Taxation in software industry has been debated for many decades (OECD, 2015). The problem with taxing the final result of software engineering, the software product or service, for example, has shown to be challenging to tackle and is still subject to ongoing and broad discussion (Olbert and Spengel, 2017). This debate has also reached software engineering, the construction of the software product; at first in the context of inner source, use of open source software development practices and the establishment of open source-like communities within an organization (Capraro, 2020). To address the taxation challenges in inner source software development, Buchner and Riehle proposed applying a so-called cost-plus approach (Buchner and Riehle, 2022). However, Treidler et al. showed that such a cost-plus approach is not applicable since inner source does not reflect the function and risk profile, the starting point for any transfer pricing discussion, for a cost-plus approach (Treidler et al., 2024a).
@@
Dorner et al. identified the use of software components owned by teams or individuals who represent separate geographically distributed subsidiaries of the same enterprise as a type of licensing—a key intangible in software engineering that is taxable (Dorner et al., 2024). We will explain the concerns of tax authorities on software architecture in Section 4.1, in laymen’s terms and with a simplified example of a multinational enterprise developing software, how tax authorities become a stakeholder in a globally distributed software architecture.
The very recent litigation from 2023 between Microsoft and the US tax authorities (IRS), in which the IRS alleges that Microsoft owes an additional \$28.9 billion in tax from 2004 to 2013, plus penalties and interest, demonstrates the importance of tax compliance for multinational enterprises developing software. For the interested reader, we recommend our analysis of the Microsoft case from a legal perspective (Treidler et al., 2024b).
3 Study Design
Our study consists of two parts as Figure 1 highlights:
  1. 1.
    An exploratory part, in which we develop a viewpoint to enable software-developing multinational enterprises to report their implicit license structure derived from the author team’s expertise and understanding of the topic and construct a view of a multinational enterprise’s software architecture
  2. 2.
    An confirmatory part, in which we evaluate the resulting software architecture description for a large-scale software architecture of a multinational enterprise with a panel of tax experts.
[…]
Figure 1: An overview of the structure of our study: (1) creating a software architecture description with a viewpoint derived from the concerns of tax authorities and a view of an empirical software architecture using the viewpoint (exploratory) and (2) evaluating the software architecture description with a panel of tax experts (confirmatory).
As the main contribution of this study, we focus on the evaluation (confirmatory) part of our research on tax compliance in the context of software architecture. To evaluate the software architecture description for its intended purpose, we conducted semi-structured interviews with four tax experts which makes our study a judgment study. In a judgment study, researchers gather empirical data from a group of participants who are asked to judge or rate behaviors, to respond to a request or “stimulus” offered by a researcher, or to discuss a given topic of interest (Stol and Fitzgerald, 2018). Stol and Fitzgerald (2018) compare judgment studies to a courtroom: a panel of participants (the jury) is carefully and systematically selected. In a courtroom, evidence is presented (a stimulus) and eventually the jury returns a verdict. The setting itself (i.e., the courtroom) is only manipulated to the extent that it aims to be neutral and not distract the participants from the matter at hand (i.e., the case). The goal of a judgment study is to seek generalizability over the responses, rather than generalizability to a population of actors (Stol and Fitzgerald, 2018).
Therefore, we systematically sampled our experts from both opposing perspectives in a tax audit:
  • the company perspective as the provider represented by a tax advisor
  • the tax authority perspective as the recipient of such documentation represented by a tax auditor
We selected the participants from our professional network. None of the participants were compensated for their participation. Additionally, none of the participants are currently working or have previously worked with any members of the author team, neither in an academic nor in a professional context. Table 1 provides an overview of our interviewees and their backgrounds.
Table 1: An overview of the interviewees and their backgrounds.
@@ 3 Study Design
A Tax auditor Former tax auditor, majoring in international taxation, OECD advisor for transfer pricing from 2016 to 2018, currently not involved in transfer pricing
@@
K Tax advisor Transfer pricing advisor, previously transfer pricing expert at one of the Big Four accounting firms for 20 years
@@
M Tax auditor Tax auditor and key expert on transfer pricing for other tax auditors for more than eight years
@@
R Tax advisor Managing director at a transfer price consulting firm; tax inspector before, 35 years of experience in international taxation
The stimulus for the panel of four tax experts is a software architecture description of the implicit license structure of a globally distributed software architecture of a multinational enterprise.
We ensured privacy and confidentiality for all actors (i.e., the involved companies, tax authorities, and tax advisors) by conducting the interviews separately and removing all company-specific information. We also altered the geographical distribution of the subsidiaries to make our company not identifiable in any sense on the one hand but also to give the reader a better impression of globally distributed software architecture on the other hand. The interview guidelines are publicly available [^1] https://zenodo.org/records/10214942.
We conducted two pilot interviews with the author team and a third tax auditor (who was not part of the panel) to address didactical challenges since the panel has no software engineering but a taxation background. A pair of a co-author with a taxation background and a co-author with a software engineering background conducted the interviews with the tax experts virtually between 2023-10-24 and 2023-11-03. We recorded all interview sessions. The interviews were held in English and German.
@@
As our interview guidelines show, large parts of the interviews were dedicated to providing contextual software engineering understanding for the tax experts. Therefore, we translated and paraphrased only the answers from the interviewees. We analyzed the interviews in a team of two: While one researcher of the interview team analyzed the recorded interviews using thematic analysis (Clarke and Braun, 2017), the second one double-checked findings and mappings in a separate session. To avoid translation mistakes and misunderstandings, the final version of the paper was sent to each interviewee, asking for confirmation.
4 A Software Architecture Description for Tax Compliance
@@
In the following subsections, we first provide a gentle introduction to international taxation for software engineers in Section 4.1. We then derive based on our interdisciplinary experiences and opinions the viewpoint, which is the set of conventions for creating, interpreting, and using an architecture view to frame the concern of tax authorities (Section 4.2). In Section 4.3, we apply then the viewpoint to a globally distributed software architecture of a multinational enterprise to be able to evaluate a full software architecture description with the panel of tax experts in Section 5.
4.1 International Taxation Standards as Concern and Tax Authorities as Stakeholder
@@
Consider the fictional devnullsoft Group, a multinational enterprise that develops and sells a software product. The software is split up into 18 components. Each component has an owning team that is responsible and accountable for its components.
The teams are assigned to one of the legal entities of the enterprise in three countries depending on the location of the team members: devnullsoft GmbH in Germany, devnullsoft AB in Sweden, and devnullsoft Ltd. in the UK. In Figure 2, we provide a schematic overview of a simplified globally distributed software architecture with owning teams assigned to three different subsidiaries in Germany, Sweden, and the UK. Using code components across different jurisdictions is taxable.
@@
Figure 2: A schematic overview of a globally distributed software architecture developed by the fictional devnullsoft Group, a multinational software enterprise with three subsidiaries developing software. Using code components across different jurisdictions is a form of licensing and, thus, taxable.
@@
All three questions must be answered by a software architecture description to ensure tax compliance for multinational enterprises.
4.2 Architecture Viewpoint
@@
4.2.1 Defining component and component owner
@@
4.2.2 Decomposing the component structure of the software product
@@
4.2.3 Identifying component owners
@@
4.2.4 Identifying jurisdictions of component owners